### v1

# Sub-keV dark matter can strongly ionize molecular clouds

Abstract
We show that the ionization of dense molecular clouds can be used to set strong constraints on dark matter models producing UV/X-ray photons in their annihilation or decay. We place robust and competitive constraints on various dark matter models, such as axion-like particles, scalars and sterile neutrinos, for masses between eV and keV, and project forecasts to illustrate the potential of this target. We discuss how these constraints can be significantly improved by considering a more refined sample of molecular clouds near the Galactic Center and above the Galactic plane, a detailed modeling of the cosmic-ray ionization contribution and, potentially, a more refined analysis of the gas density in clouds through dust extinction maps. Thus, ionization of molecular clouds emerges as one of the most powerful tools for probing sub-keV dark matter.
I Introduction
Despite compelling evidence for the existence of dark matter (DM) through its gravitational effects, the fundamental nature of DM remains largely unknown. If DM is indeed a new fundamental particle, it is crucial to explore its possible interactions beyond gravity; see Ref. Cirelli et al. (2024) for a comprehensive review. Numerous studies have placed stringent constraints on heavy fermionic (WIMP-like) DM with masses in the GeV range and above Aalbers et al. (2024); Bo et al. (2025); Aprile et al. (2025); Agnese et al. (2018); Agnes et al. (2023); Ajaj et al. (2019); Akerib et al. (2019). Significant efforts are currently underway to probe lighter DM candidates through both direct and indirect detection strategies Antel et al. (2023). These approaches aim to uncover signals from particles light enough that escape the traditional WIMP-focused searches. A wide array of theoretical models extending the Standard Model (SM) suggest the existence of light, weakly interacting particles that could account for some or all of the present-day DM density. These extensions often emerge in attempts to address unresolved issues in particle physics, such as the smallness of neutrino masses or the strong-CP problem. Among the most extensively studied candidates in this mass range there is the QCD axion Peccei (2008), a pseudoscalar particle originally proposed to solve the strong-CP problem, and more general axion-like particles (ALPs) Choi et al. (2021), which appear in many string-theoretic frameworks and span a broad range of masses and couplings Gendler et al. (2024). Significant effort is devoted to sterile neutrinos Dasgupta and Kopp (2021), which are singlets under the SM gauge group but may mix with active neutrinos and contribute to the DM and neutrino phenomenology, and dark photons Fabbrichesi et al. (2020), hypothetical gauge bosons associated with a hidden symmetry that can kinetically mix with the SM photon. The search for these light DM candidates employs a diverse set of experimental techniques. Direct detection experiments attempt to measure the scattering or absorption of such particles in terrestrial detectors Lanfranchi et al. (2021), while indirect detection methods aim to observe their decay or annihilation products in astrophysical environments Raffelt (1996); Caputo and Raffelt (2024); Jacobsen et al. (2023); Foster et al. (2022); Carenza et al. (2025).
In this work, we follow this latter path and propose a new method to probe the coupling of DM with masses at the keV scale and below. In particular, we show that the annihilation/decay of such low-mass DM particles into photons from ultraviolet (UV) to X-ray energies is able to produce significant ionization inside molecular clouds (MCs), reaching levels of ionization even above the ionization that cosmic rays (CRs) are expected to cause. Interactions between DM, or its annihilation/decay products, and the visible sector could ionize molecular hydrogen, , which is abundant within MCs. For example, DM decay into photons can lead to ionization when the DM mass satisfies eV, since the ionization energy of molecular hydrogen is approximately eV, producing high levels of ionization when these photons are absorbed inside the MC. Similarly, a coupling between DM and electrons can cause ionization through scattering processes. In dense molecular media, and, in particular, in dense MCs, UV photons are quickly absorbed, delivering all of their energy into the medium and ionizing the gas. Since DM would be nearly homogeneously distributed inside these clouds (extending over scale order of parsecs), its annihilation/decay would lead to an intense, homogeneous, ionization inside such systems. This potential DM-induced ionization contributes alongside standard ionization sources, such as stellar UV radiation, low-energy CRs, and X-rays emitted by stars within the MC Krolik and Kallman (1983); Norman and Silk (1980). It is possible to set strong constraints on the DM lifetime, annihilation rate, or on the coupling to photons when DM generates a ionization that exceeds the standard astrophysical sources.
Previous literature already investigated the heating of gas Bhoonah et al. (2018); Wadekar and Farrar (2021); Alvey et al. (2023) or the ionization Blanco et al. (2024); Balan et al. (2024) induced by the direct interaction of DM particles with the nuclei in gas systems, such as rich-gas dwarf galaxies, MCs or the early universe, through the cm line Evoli et al. (2014) for example. However, the effect of particle emissions from DM in the ionization of MCs has not been investigated before, making this the first effort to illustrate the potential of such clouds to indirectly probe light DM. As we show, on astrophysical scales, MCs act as large-volume detectors with ultra-low thresholds for ionization effects.
This paper is organized as follows: In section II we overview the expected main drivers of ionization in MCs and briefly outline the measurements of the ionization rate. Then, we discuss the targets that we are going to use in our study and identify optimal MCs to constraint DM. In Sec. III we detail our calculation of the ionization caused by photons produced from DM decay/annihilation. Then, we describe our results is Sec. IV and summarize our main findings in Sec. V, where we discuss possible ways of optimizing our current constraints.
II Ionization of dense molecular clouds
While MCs with low column densities111The column density is defined as the integral of the density of a system over our line of sight, , N cm-2, called diffuse MCs, are expected to be highly affected by external UV radiation penetrating the cloud, above this column density threshold, corresponding to dense MCs Snow and McCall (2006), the UV radiation is strongly attenuated, and low-energy CRs are expected to be the dominant ionizing agent McKee (1989). In fact, CRs are expected to drive the chemical complexity observed in neutral gas systems and play a major role in the regulation of star formation. Modeling the contribution of CRs to the ionization of MCs is an intricate problem. Early studies estimated that the ionization rate per atom in clouds composed mainly of atomic hydrogen is within the range of a few Spitzer and Tomasko (1968); Indriolo and McCall (2012); Phan et al. (2023); Indriolo et al. (2015). This broad uncertainty reflects poor knowledge of the CR spectra at energies much below MeV. In particular, the ionization cross section for atomic hydrogen peaks at approximately eV for CR electrons and around keV for CR protons Padovani et al. (2009). In order to determine with more accuracy the expected ionization rates, a better understanding of the low-energy spectra of CR electrons, protons, and heavier nuclei is required. This is an extremely challenging task. The simplest approach consists of measuring the low-energy CR spectrum reaching Earth and extrapolating it to energies below the CR observations. Unfortunately, local measurements of the CR spectra at low energies are affected by solar modulation and the structure of the interplanetary magnetic field. The Voyager I and II spacecrafts have provided valuable data, measuring the electron spectrum down to a few MeV Cummings et al. (2016); Stone et al. (2019) beyond the heliopause, although these observations are still far from the energy range most relevant for ionization. An alternative to probing these low energies involves observations of the Galactic synchrotron background. This method is limited by significant uncertainties in modeling the interstellar magnetic field Padovani and Galli (2018); Orlando (2018); Ginzburg and Syrovatskii (1965) and it would hardly probe energies of tens of eV without involving a low-energy extrapolation of the CR spectrum. Thus, modeling the ionization of dense MCs is subject to significant uncertainties.
Experimentally, measuring the ionization of MCs is of interest. A particularly effective method to infer the CR ionization rate involves measuring the intensities of emission lines from molecular tracers, whose abundances are closely tied to the free electron population. The choice of tracer strongly depends on the gas density and chemical composition. In diffuse regions of MCs, is a favorable tracer due to its simple and well-understood chemistry Indriolo and McCall (2013). In denser environments, such as MC cores and proto-stellar clusters, commonly used tracers include HCO+, DCO+, and CO Caselli et al. (1998). At present, measurements of the ionization rate in MCs seem challenging to explain by CR ionization. First, if CRs are the dominant ionization agent of MCs across the Galaxy, one expects that the ionization rate decreases for MCs with larger column densities. Although there are observations that support this trend, no clear correlation is observed Gabici et al. (2019); Gabici (2022). Second, the levels of ionization from CRs are expected to be lower than that measured in several clouds Indriolo et al. (2009); Phan et al. (2018). In this context, the high ionization rate observed in the Central Molecular Zone have been found to be incompatible with what the ionization that CRs can provide Ravikularaman et al. (2025), and MeV DM has been proposed as a potential explanation of this anomalous ionization De la Torre Luque et al. (2025). In addition, there are hints from gamma-ray observations suggesting that CRs do not actually penetrate dense MCs Yang et al. (2023).
As opposed to standard ionization mechanisms, UV photon produced by DM interactions would ionize directly within the cloud, providing a constant additional ionization to that provided by CRs. In this work, we explore the potential of dark, or dense, clouds, which are observed with temperatures near K, sizes from a below the parsec to tens of parsecs and masses up to M⊙ Bergin and Tafalla (2007), to probe the emission of UV/X-ray photons from sub-keV DM. In order to show how promising are such clouds as targets for DM searches, we will derive constraints from a local dark cloud as a pessimistic (and robust) case, because the modeling of the DM density does not need assuming any DM density distribution and will therefore be very robust. Then, we will evaluate constraints from a cloud closer to the Galactic Center (GC), where DM is expected to be concentrated, and we will assume a prototypical Navarro-Frenk-White (NFW) Navarro et al. (1996) distribution. This is the standard assumption made in most of the indirect DM searches, and allows for a fair estimation with previous constraints. As we will explain below, we follow a very conservative approach in both cases, and better results can be obtained through a more dedicated search of optimal clouds that is left for future work.
One of the local clouds that has received more attention is L1551, which is located pc away from the Earth in the direction of the Taurus constellation and has a size of pc. This cloud presents a low ionization rate, with de Boisanger et al. (1996). We will derive constraints at confidence level by requiring the DM-induced ionization to match the observed ionization at the level. Moreover, we note that we are not considering here the standard ionization that would be produced by CRs or UV/X-rays, which makes our estimations extremely conservative. We will discuss the expected improvement when CR ionization is modeled.
We stress that L1551 is not the local cloud with the lowest ionization rate. For example, the Taurus molecular cloud (TMC-1) is a dark cloud in the constellations of Taurus and Auriga. For this cloud, Ref. Flower et al. (2007) found that provided the best overall agreement with the observations. Using this cloud would improve our limits by a factor on the DM annihilation/decay rate. However, the lack of a robust estimation of the uncertainties of this measurement makes us choose L1551 as reference case. Another interesting example is Barnard 68, which was found to be compatible with an ionization rate of Maret and Bergin (2007). Unfortunately, there is no precise estimation of the uncertainty of these measurements that allow us to use them reliably, but it must be noted that using these clouds could lead to up to an order of magnitude stronger DM constraints than our reference case. Lower values of ionization can be found in the local environment, but we leave a dedicated survey for a future work.
Interestingly, even lower values of ionization have been found in infrared dark clouds (IRDC), which are very cold (T K) and dense clouds, and can be ideal targets for DM searches. Among known IRDCs, one that has been very well studied recently is G28.37+00.07, often referred to as the DRAGON cloud, which is located around kpc away from the Galactic center (in the so-called -kpc ring) Simon et al. (2006). This cloud has a size around L pc, a density of n cm-3 and a column density N cm-2. Ref. Entekhabi et al. (2022) studied the ionization rate and other gas properties across different sub-regions of this cloud, finding that the ionization rate in some sub-regions was exceptionally low. Indeed, the authors of this paper chose some of these sub-regions because they were expected to gather no star formation activity. Since the ionization from photons produced by DM is expected to be uniform throughout the cloud, due to negligible variations in DM density on scales of a few parsecs, constraints on DM-induced ionization can be derived from measurements in individual zones within the cloud. In particular the region P8 of that study was found to have ionization values between s-1 and s-1 in their fiducial model. We set an upper limit on the ionization of this cloud by taking twice the highest value found in this analysis, i.e. s-1. This will correspond with our fiducial constraint.
While the modeling of DM ionization in local clouds would not require to assume any DM distribution (i.e. leading to really robust bounds), we select a cloud located quite close to the GC to demonstrate the potential of such dark clouds to probe sub-keV DM. The advantage is the higher DM density in this region, with the disadvantage of introducing uncertainties related to the DM distribution profile. For concreteness, we choose G1.4-1.8+87, with a size of around pc, and located around pc away from GC, above the Galactic plane Bhoonah et al. (2018). This cloud was found to be compatible with an ionization rate as low as Bhoonah et al. (2019), however, Ref. Farrar et al. (2020) casted serious doubts on the gas temperature used in this analysis. We will use this cloud as an example of the optimistic case of MC with low ionization rate located close to the Galactic Center. However, this should not be interpreted as a strict constraint, but rather as a forecast to illustrate the potential of similar clouds once more robust analyses are performed. To derive this limit, we choose a value of ionization of and will assume a prototypical NFW DM profile, which is the standard assumption in this kind of searches. We will call this an optimistic constraint, that will be shown in a different way than the fiducial and conservative cases, for clarity. As in the case of local MCs, more research is needed to identify more clouds close to the GC. IRDCs located close to the GC but above the Galactic plane are probably the best target to constraint the ionization caused by sub-keV DM, since the impact of CR ionization will be significantly reduced.
III Ionization of UV/X-ray photons from keV DM
The ionization rate , for a MC located at x caused by photons with energy , can be calculated as Padovani et al. (2009); Phan et al. (2023)
where is the photon flux in a given location of the sky, is the photoionization cross section Padovani et al. (2009) and is the number of secondary ionizations per primary ionization Yan et al. (2001); Dogiel et al. (2014)
Here, eV is the ionization potential of H2, is the energy of the electron produced from the photoionization induced by an incoming photon with energy and eV is the average energy loss of a non-thermal electron per ionization of an H2 molecule in the interstellar medium Dalgarno et al. (1999). The integral in Eq. 1 extends from a minimum energy of to the maximum photon energy produced in the DM annihilation/decay, i.e. the mass of the DM particle, mχ, in the case of annihilating DM and mχ/2 in the case of decaying DM. Here, we are neglecting the Compton ionization process, which is important only above tens of keV Yan et al. (2001). This simplification is motivated by the fact that at such energies, our limits are not competitive with previous constraints.
The photon flux at a given position can be expressed as
where the factor accounts for the production of two photons, LMC is of the size of the MCs (LMC) Gabici (2022) calculated for a constant average n and is the source term for the injection of photons
Here, and are the DM thermally averaged annihilation cross section and the DM decay rate, respectively. The DM energy density at the position x is and is the -ray yield per DM interaction, that is a Dirac delta peaking at mχ, in the case of annihilation, or at mχ/2 in the case of decay.
In Fig. 2 we show the quantity J for several DM masses in a local MC like L1551. As expected, the lower the DM mass, the higher is J and, thus, the ionization rate. This behavior is due to the higher DM number density and the higher photoionization cross sections, which peaks around eV Yan et al. (2001). This can be also seen from Fig. 2, where we show the H2 ionization rate caused by annihilating (dot-dashed line) and decaying (solid lines) DM in a local MC. The purple dashed line shows the level of ionization observed for L1551. In this work we will constrain the annihilation rate into photons defined as , where is the total DM annihilation rate and is the branching ratio for photons in the final state. Similarly, we will always refer to DM decay into photons.
For the local cloud L1551, we set the DM density to be the local DM density, which is GeV/cm3 Benito et al. (2021) at the Earth position kpc from the Galactic center Evans et al. (2019). We stress that this makes our estimation independent on the DM density profile, which is one of the main uncertainties in DM indirect searches. In fact, given that the considered cloud is located in the Solar System neighborhood, this assumption leads to a quite conservative result. For the case of the DRAGON cloud, which corresponds to our fiducial constraint, we assume the NFW profile Navarro et al. (1996), as it is the most usual assumption used for DM searches. Considering constraints on the DM lifetime , using the NFW profile leads to a factor stronger bound than assuming the local DM density at the cloud position, while for annihilation () the difference is close to a factor . We stress that cuspier DM profiles, like the ones favored by the GC excess Ackermann et al. (2017), are commonly used as optimistic scenarios, but we will not consider them. We also adopt a NFW profile for our optimistic forecast with G1.4-1.8+87. We also note that the local DM density value that we have taken is in the lower bound given in Ref. Benito et al. (2021), where the estimated local DM density is GeV/cm3 at 2 level.
IV Results
In Fig. 3, we report our model-independent constraints on DM two-photon decay (left panel) and annihilation (right panel). In this energy range, the strongest constraints are from CMB spectral distortions Bolliet et al. (2021) and heating of gas-reached dwarf galaxies (particularly, the Leo T dwarf) Wadekar and Wang (2022). The latter constraints not only requires the adoption of a model for the DM distribution, gas temperatures and the density distributions of molecular and atomic gas phases as well as dust grains (which are uncertainties that can similarly affect the measurement of the ionization rate of the DRAGON cloud), but it also requires modeling the background heating from CRs and UV/X-ray. Thus, different assumptions imply different uncertainties that are involved in these constraints. Here, we show that even in the most pessimistic case of a local MC such as L1551, with a low DM density, the constraints at tens of eV are competitive with the existent bounds Wadekar and Wang (2022); Bolliet et al. (2021); Audren et al. (2014); Poulin et al. (2016). Notably, our fiducial bound (the DRAGON cloud), becomes stronger than CMB constraints in the mass range from eV to almost keV, and it is also stronger than the LeoT gas heating constraints up to a few hundreds eV. As commented above, the assumption of a NFW profile for the DRAGON cloud makes our constraint only a factor of stronger than assuming only the local DM density at the cloud position, and around a factor of for the annihilation rate constraint. Taking a Burkert profile Burkert (1995), which is often taken as the most pessimistic scenario in Galactic DM searches, these limits are expected to be reduced by a factor for decay and for annihilation. We remind the reader, that even cuspier profiles than NFW are equally likely, and those would result in an improvement of this limit. Notably, for the optimistic forecast with G1.4-1.8+87 we get a constraint that can only be improved by direct X-ray observations, such as those from XMM-Newton Gewering-Peine et al. (2017); Foster et al. (2021) above the keV. We tested another interesting candidate, the cloud G357.8-4.7-55, which is around kpc Bhoonah et al. (2018) far from the Galactic Center and is also compatible with low ionization rates. Assuming a NFW, we get a constraint that is more than one order of magnitude better than the one from L1551 for the lifetime constraint.
We have also estimated what could be the effect of adding the ionization from CRs in L1551. Since CRs are expected to be the ionization agents in the interstellar medium, their effect will be dominant. In fact, previous predictions expect the ionization of L1551 to be totally explained by CRs Phan et al. (2023). To remain conservative in our estimation, we adopt the CR ionization calculated in Ref. Phan et al. (2023), pointing towards an ionization rate between s-1 and s-1 in the diffusive regime. This will be the contribution to the ionization of L1551 explained by standard physics. We adopt the most conservative value of this estimation, s-1 in our calculation. We find that adding this contribution leads to around a factor of better constraints below keV for both decay and annihilation. This rough approach illustrates the improvement expected in the derived limits when accounting for the background ionization of CRs. A similar estimation could be made for clouds that are located farther away from the Solar System, but this would require a precise modeling of the distribution of low-energy CRs, and the uncertainties would be very important.
These model-independent constraints for DM producing two photons can be directly expressed as constraints in the coupling of different feebly interacting particles, such as ALPs, scalars or sterile neutrinos. We calculate our constraint on the coupling of ALPs with photons as Carenza et al. (2025)
where is the model independent limit on the DM lifetime and ma is the ALP mass, which is from 10 eV to 10 keV in our analysis. The emergence of ALPs in the eV-keV mass range might be unrelated to solving the strong CP problem, and for this reason we refer to them as ALPs and not QCD axions Peccei (2008); Choi et al. (2021); Gendler et al. (2024); Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978). Several QCD axion and ALP models predict viable DM in the eV–keV range, motivating searches in this regime Di Luzio et al. (2021); Giannotti (2005); Di Luzio et al. (2020); Rubakov (1997); Gavela et al. (2024); de Giorgi and Ramos (2025); O’Hare et al. (2022); Gelmini et al. (2021); Ayala et al. (2014); Carenza et al. (2020); Lucente et al. (2022).
In Fig. 4 we report our bounds for the coupling of ALPs with photons, gaγ, for keV mass ALPs, compared to previous constraints. We highlight here that, even in the most pessimistic case of L1551, these constraints are competitive with constraints from reionization Cadamuro and Redondo (2012) and extend to higher masses. For our fiducial constraint, from the DRAGON cloud, we significantly improve on the CMB and LeoT constraints up to several hundreds of eV. We also point out that constraints based on observations of Leo T are influenced by different sources of uncertainty in their model of the gas distribution, their temperature and its DM halo Wadekar and Wang (2022), that are expected to be similar to those affecting the ionization rate measurement for the DRAGON cloud. The halo model used to derive their constraint is discussed in Ref. Faerman et al. (2013), whose uncertainties can mildly affect the resulting constraint. However, a more important issue is connected to the possible breaking of the hydrostatic equilibrium assumption, fundamental for the model of Ref. Faerman et al. (2013). Indeed, even though observations do not point to a rotation of the HI gas in Leo T, it is possible that the gas is rotating in a direction aligned with our view, which would conceal its motion. However, Ref. Wadekar and Wang (2022) claimed that these uncertainties are under control and the failure of hydrostatic equilibrium might even lead to a stronger bound. Moreover, our optimistic forecast for G1.4-1.8+87, that represents what we expect from a low ionization cloud near the Galactic center, illustrates that MC observations could be the leading constraints in the sub-keV range for ALPs and similar particles.
We also report the constraint for scalar particles as DM, that have mixing angle with the SM Higgs Dev et al. (2021); Nguyen et al. (2024). We calculate this mixing angle from the scalar dark matter decay to two photons lifetime as
where is the fine structure constant, GeV is the electroweak vacuum expectation value and mS is the mass of the scalar particle. A comparison of our constraints with previous ones, for scalar DM, is shown in Fig. 5. As for the case of ALPs, we are able to improve on existent constraints, such as those from CMB spectral distortions Bolliet et al. (2021), LeoT gas heating Wadekar and Farrar (2021) or the model independent constraint on decaying DM (dDM) derived in Refs. Audren et al. (2014); Poulin et al. (2016), and our results are complementary with X-ray constraints Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021), constraints from the CMB anisotropies Slatyer (2016) and stellar constraints Dev et al. (2021); Balaji et al. (2022). Interestingly, MCs can probe a region of the parameter space that cannot be currently tested from arguments of the stars dynamics.
We also compute our limit for sterile neutrinos, which have a decay of the form . The expression used for our calculation is Boyarsky et al. (2019); Bezrukov et al. (2010)
where m is the sterile neutrino mass and GF is the Fermi constant that represents the strength of the weak force. In this case, instead of producing two photons in the decay, only one is produced. Therefore, the value adopted is a factor of 2 lower than the one shown in Fig. 3. These bounds are shown in Fig. 6, compared to previous ones. As one can see, the parameter space that is covered by MCs is already well tested by other mechanisms, specially the Lyman-alpha forest Bringmann et al. (2023). Also the Tremaine-Gunn constraint Tremaine and Gunn (1979) rules out sterile neutrinos with masses below keV. We report the optimistic constraint derived from G1.4-1.8+87, which may become relevant again if model-dependent approaches reopen this region of parameter space Dev et al. (2025).
Finally, we have also calculated these constraints for the case of vector bosons decaying into three photons. In this case, the spectrum of photons is given by a continuum spectrum, instead of a line signal. The photon flux at a given position can be expressed as Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b)
where . Inserting this term in Eq. 4, we can compute Jχ→3γ as in Eq. 3, for three-photon final state. The result is shown in Fig. 7, upper panel. A trend similar to the monochromatic injection appears, with a larger ionization at lower DM masses. In Fig. 7 (lower panel) we show the limit on the inverse decay rate of three-photon final state injection, for the case of L1551 and G1.4-1.8+87.
We derive constraints on well-motivated vector dark matter models, such as the dark photon and the vector boson, using decay rate limits of the vector-to-three-photon final state. Both models arise from extensions of the Standard Model by anomaly-free gauge symmetries. The dark photon interacts with Standard Model particles via kinetic mixing with the photon field Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021), while the vector boson couples directly to the Standard Model through a flavor-universal coupling Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024). These symmetry extensions are extensively studied in the contexts of string theory, string compactification, and vector dark matter production via quantum fluctuations during inflation Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b).
Figure 8 presents the optimistic G1.4-1.8+87 constraints on vector dark matter models. We compute the decay width using the Euler-Heisenberg (EH) approximation, which is valid in the keV mass range McDermott et al. (2018); Linden et al. (2024b). The top panel shows the constraint on kinetic mixing , compared to limits from the universe’s age (4.6 Gyr), x-ray observations by INTEGRAL Linden et al. (2024b), and direct detection bounds from XENON Aprile et al. (2019). For vector dark matter, we derive constraints on the flavor-universal coupling , assuming the decay proceeds dominantly via electron loops under the EH approximation. We compare this result with the INTEGRAL limit on the same three-photon final state Linden et al. (2024b) and with theoretical constraints based on neutrino decay channels and the universe’s lifetime Chun and Yun (2022).
For the dark photon, the G1.4-1.8+87 constraint surpasses the theoretical bound, as the three-photon decay dominates in this mass range. However, this region is already excluded by direct detection experiments that consider dark photon absorption. In contrast, our constraints lie above the bound set by the universe’s lifetime, assuming decay into neutrino pairs—the dominant channel for the vector boson. This behavior extends to other lepton-flavor-dependent vector dark matter models such as . Overall, our results represent conservative limits on these classes of vector dark matter models.
V Discussion and conclusions
In this paper, we investigate the ionization of neutral gas produced in dense MCs by decaying/annihilating DM. In particular, sub-keV DM could generate UV/X-ray photons that are able to produce a high and homogeneous ionization across the cloud since they are absorbed efficiently within such systems. In fact, we show that the ionization produced by the photon emission from sub-keV DM saturates the ionization that is observed in different MCs for decay lifetimes, or annihilation rates, that are not ruled out by any other constraint. Therefore, we test for the first time MCs as a target for indirect DM searches, demonstrating that they allow us to improve current constraints for decaying DM particles such as ALPs, dark photons, scalar particles or sterile neutrinos for masses from a few tens of eV to slightly below keV.
Interestingly, The ionization of dense MCs is expected to be primarily driven by CRs, as UV light cannot effectively penetrate these regions. However, sub-keV unstable DM particles could be producing UV radiation directly inside the MC, leading to a persistent background of ionization, potentially influencing both the star formation rates and the chemical evolution within the clouds. Notably, a recent study suggests that DM particles with MeV-scale masses might account for the anomalously high ionization rates observed in the Galactic Center De la Torre Luque et al. (2025).
We begin by identifying local MCs as a particularly robust probe of this kind of DM. This is because estimating their ionization rate requires only knowledge of the local DM density — without the need to assume a specific DM distribution model. Using the local MC L1551, which is not the cloud with the lowest observed ionization rate, we derive constraints that are similar, or even better, than the previous leading constraints. Many of the existing bounds rely on assumptions that can affect their reliability and expose them to significant uncertainties, as opposed to the more robust MC bounds that we explored. Then, we examine the case of G28.37+00.07, or the DRAGON cloud, which has been studied previously in detail and shows regions of it that are exceptionally quiescent in terms of their star formation activity, with ionization rates compatible with s-1. This constitutes what we call our fiducial constraint, and allows us improve previous constraints from CMB and LeoT gas heating up to several hundreds of eV. Finally, we use the case of G1.4-1.8+87, an MC that is pc away from the Galactic Center, to show the potential of using clouds near the Galactic Center. For this case, we take the values obtained in Ref. Bhoonah et al. (2019) to illustrate that, assuming a NFW DM density profile, we reach the strongest DM bounds to date in a mass range from a few tens of eV to the keV. However, this must be taken as an optimistic forecast since the characterization of the temperature of that could was put in doubt in Ref. Farrar et al. (2020). We highlight that clouds that are closer to the Galactic Center can be found, therefore providing even stronger DM bounds. Finally, we highlight that the most promising targets for indirectly searching for sub-keV DM decays or annihilations into photons are IRDCs located near the Galactic Center but slightly above the Galactic plane. These regions are expected to be less affected by CR ionization, which is sourced in the Galactic disk. As a result, such clouds are likely to exhibit very low ionization levels, making any DM-induced effects more distinguishable. Although a more comprehensive survey of MCs will further refine the constraints presented here, our results already demonstrate the strong potential of this approach for probing different types of light DM particles. We also note that this strategy can be also applied for MCs in nearby galaxies.
Among the expected future refinements, we list a few below.
i) A more exhaustive survey of the measurements of dark MCs, dedicated to identify clouds with lower ionization fractions. These measurements are often complex, and they are usually focused on measuring the ionization around clouds with regions of star formation. However, we show that dark clouds constitute a very interesting target for DM searches, hoping to increase the interest towards targeting dark MCs for observations. In addition, if dark clouds systems can be found closer to the GC, our limits will be highly improved. In this regard, there are already extensive surveys of such clouds, like the analyses from Ref. Simon et al. (2006). A dedicated exploration and analysis could yield improvements of up to a few orders of magnitude in the derived DM constraints.
ii) A recent paper Obolentseva et al. (2024) showed that previous measurements of the ionization rate from several dense MCs are biased, due to an inaccurate evaluation of their density. Using dust extinction maps to evaluate their density leads to an ionization rate that is lower by an average factor of . Analyzing interesting targets using this procedure could lead to constraints that are an order of magnitude better than the current ones.
iii) Probably the most obvious improvement would be that of modeling the ionization produced by CRs. Provided that CRs are expected to be the main ionization agent of dense MCs, including their contribution must certainly improve our constraints. Unfortunately, this estimation is not simple and requires an accurate knowledge of the distribution of CRs in different regions of the Galaxy and at very low energies, where our models are not robust enough to produce a reliable constraint. However, reasonable estimations can be made for local MCs, and improvements of a factor of a few in the constraints can be already obtained with conservative assumptions, as discussed above.
As a final point, we also note that X-ray telescopes are very promising for constraining DM above a few hundreds eV. Especially eROSITA could place constraints above keV that can beat all previous constraints Dekker et al. (2021), once they have enough observational time. However, the the current constraints from eROSITA or Chandra below keV are even weaker than those that we obtain and forecast in this work.
In conclusion, we have shown the potential impact of sub-keV DM decaying/annihilating into photons in dense gas regions, which can be specially relevant for probing ALPs or scalar DM particles. Our results demonstrate that astrochemical observations in MCs can serve as a novel and powerful probe for indirect searches of sub-keV DM. This approach offers particularly robust constraints across a range of particle physics models and has the potential to surpass existing bounds for DM masses between approximately eV and keV.
VI Acknowledgements
We thank Torsten Bringmann for fruitful discussion. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. PC and TTQN are supported by the Swedish Research Council under contract 2022-04283. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).