### v2

# Sub-keV dark matter can strongly ionize molecular clouds

Abstract
The ionization of dense molecular clouds offers a powerful and robust probe of dark matter models that produce UV or X-ray photons via annihilation or decay. Using this novel observable, we derive competitive constraints on dark matter in the mass range from 30 eV to 1 keV. In particular, we set the strongest existing limits on axion-like dark matter in the 30–100 eV range. We also present sensitivity projections demonstrating the significant potential of this approach. These constraints can be substantially improved with targeted observations of molecular clouds near the Galactic Center and above the Galactic plane as well as with a more accurate modeling of cosmic-ray ionization and refined gas density estimates from dust extinction maps. Our results establish molecular cloud ionization as one of the most sensitive tools for probing sub-keV dark matter.
Introduction — Uncovering the fundamental nature of dark matter (DM) remains one of the most pressing challenges in physics. Therefore, it is essential to investigate possible non-gravitational interactions of DM with the visible sector Cirelli et al. (2024). Light and weakly coupled particle models are attracting significant attention Lanfranchi et al. (2021); Antel and others (2023) because of their potential in addressing other open problems, such as the strong-CP problem. The astrophysical signatures of light DM models are unique, motivating a wide range of indirect detection strategies Raffelt (1996); Caputo and Raffelt (2024); Jacobsen et al. (2023); Foster et al. (2022); Carenza et al. (2025b); Nguyen et al. (2025).
In this work we identify the ionization of dense molecular clouds (MCs) as a new and highly sensitive astrophysical probe of photons injected by light DM annihilation or decay. The photons injected by sub-keV DM have energies ranging from ultraviolet (UV) to X-ray, and they are able to produce significant ionization of molecular hydrogen, in MCs. DM-induced ionization might even exceed the ionization that cosmic rays (CRs) are expected to cause in a standard scenario Krolik and Kallman (1983); Norman and Silk (1980). Moreover, DM is a ionization source uniformly distributed across the MC, injecting ionizing radiation directly inside the dense MC. When photons are injected by DM in a dense MC, they are quickly absorbed and efficiently ionize the medium.
In the sub-keV mass range, there are many well-motivated extensions of the Standard Model (SM) predicting DM particles able, for example, to generate photons in their decay. Axion-like particles (ALPs), including the QCD axion Peccei (2008) as a special case, are generically predicted to couple weakly to photons and possibly contribute to DM Choi et al. (2021); Gendler et al. (2024). The simplest and most motivated QCD axion models Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978) can easily produce cold DM through the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983) or topological defect decay Harari and Sikivie (1987); Hagmann and Sikivie (1991); Buschmann et al. (2022); Kawasaki et al. (2015). The robust stellar bounds Ayala et al. (2014); Carenza et al. (2020); Lucente et al. (2022) usually preclude the simplest realizations of a QCD action in the eV-keV mass range. However, assuming the existence of degenerate SM copies, with the axion being the common portal in all the replicas Di Luzio et al. (2021); Giannotti (2005); Di Luzio et al. (2020), increases the QCD axion mass up to the eV-keV scale. Another possibility is that the axion couples to a copy of the QCD sector, which introduces a dynamical energy scale around GeV Rubakov (1997), increasing the axion mass at fixed photon coupling. See also Ref. Gavela et al. (2024) for models with many axions and Ref. de Giorgi and Ramos (2025) for models with extra dimensions.
The emergence of ALPs in the eV-keV mass range might be unrelated to solving the strong CP problem, and for this reason we refer to them as ALPs and not QCD axions. As a matter of fact, ALPs appear as pseudo-Nambu-Goldstone bosons of a broken global symmetry. Thus, their mass and couplings will be unrelated. The interactions between ALPs and photons lead to the production of ALPs from the thermal plasma in the early Universe. In the ALP mass range between tens of eV and keV there are stringent constraints requiring the ALP-photon coupling to be . For such a small coupling, ALPs will never reach thermal equilibrium with the visible sector, leaving the freeze-in scenario as a possibility. Non-thermal production mechanisms, such as misalignment and topological defect decay O’Hare et al. (2022); Gelmini et al. (2021), are a more appealing possibility. The misalignment mechanism, is unlikely to allow for ALP DM in the eV-keV range unless some fine-tuning on the initial conditions is involved. By contrast, depending on the details of the topological defect annihilation, ALPs in the considered mass range might compose the entirety of DM, that leave observable gravitational wave imprints Gelmini et al. (2021). Several QCD axion and ALP models predict viable DM in the eV–keV range, motivating searches in this regime.
In this Letter, we prove that the MC ionization bound can impose nearly leading constraints on axion DM from 30 eV to 200 eV, as illustrated in Fig. 1, with extremely good perspectives for future improvements. In Fig. 1, we compare existent limits with the bounds from different, well characterized, MCs (L1551 and the DRAGON cloud), and we show an optimistic forecast for a MC near the GC, G1.4-1.8+87, as described below. This exemplary DM model proves the great potential of using MCs as ultrasensitive natural detectors of DM-induced ionization.
Ionization of dense molecular clouds — While MCs with low column densities111The column density is defined as the integral of the density of a system over our line of sight, , N cm-2, called diffuse MCs, are expected to be highly affected by external UV radiation penetrating the cloud, above this column density, corresponding to dense MCs Snow and McCall (2006), the UV radiation is strongly attenuated, and low-energy CRs are expected to be the dominant ionizing agent McKee (1989). In fact, CRs are expected to drive the chemical complexity observed in neutral gas systems and play a major role in the regulation of star formation. Modeling the contribution of CRs to the ionization of MCs is an intricate problem. Early studies estimated that the ionization rate per atom in clouds composed mainly of atomic hydrogen is within the range of a few Spitzer and Tomasko (1968); Indriolo and McCall (2012); Phan et al. (2023); Indriolo et al. (2015). This broad uncertainty reflects poor knowledge of the CR spectra at energies much below MeV.
Experimentally, measuring the ionization of MCs is of interest. A particularly effective method to infer the CR ionization rate involves measuring the intensities of emission lines from molecular tracers, whose abundances are closely tied to the free electron population. The choice of tracer strongly depends on the gas density and chemical composition. In diffuse regions of MCs, is a favorable tracer due to its simple and well-understood chemistry Indriolo and McCall (2013). In denser environments, such as MC cores and proto-stellar clusters, commonly used tracers include HCO+, DCO+, and CO Caselli et al. (1998). At present, measurements of the ionization rate in MCs seem challenging to explain by CR ionization. First, if CRs are the dominant ionization agent of MCs across the Galaxy, one expects that the ionization rate decreases for MCs with larger column densities. Although there are observations that support this trend, no clear correlation is observed Gabici et al. (2019); Gabici (2022). Second, the levels of ionization from CRs are expected to be lower than that measured in several clouds Indriolo et al. (2009); Phan et al. (2018). In this context, the high ionization rate observed in the Central Molecular Zone have been found to be incompatible with what the ionization that CRs can provide Ravikularaman et al. (2025), and MeV DM has been proposed as a potential explanation of this anomalous ionization De la Torre Luque et al. (2025). In addition, there are hints from gamma-ray observations suggesting that CRs do not actually penetrate dense MCs Yang et al. (2023). As opposed to standard ionization mechanisms, UV photon produced by DM interactions would ionize directly within the cloud, providing a constant additional ionization to that provided by CRs. In this work, we explore the potential of dark, or dense, clouds, which are observed with temperatures of tens of Kelvin, sizes from a below the parsec to tens of parsecs and masses up to M⊙ Bergin and Tafalla (2007), to probe the emission of UV/X-ray photons from sub-keV DM. In order to show how promising are such clouds as targets for DM searches, we will derive constraints from a local dark cloud as the most pessimistic case, because the modeling of the DM density does not need assuming any DM density distribution and will therefore be very robust. Then, we will evaluate constraints from a cloud closer to the Galactic Center (GC), where DM is expected to be concentrated, and we will assume a prototypical Navarro-Frenk-White (NFW) Navarro et al. (1996) distribution.
One of the local clouds that has received more attention is L1551, which is located pc away from the Earth in the direction of the Taurus constellation and has a size of pc. This cloud presents a low ionization rate, with de Boisanger et al. (1996). We will derive constraints at confidence level by requiring the DM-induced ionization to match the observed ionization at the level. Moreover, we note that we are not considering here the standard ionization that would be produced by CRs or UV/X-rays, which makes our estimations extremely conservative. We will discuss the expected improvement when CR ionization is modeled.
We stress that L1551 is not the local cloud with the lowest ionization rate. For example, the Taurus molecular cloud (TMC-1) is a dark cloud in the constellations of Taurus and Auriga. For this cloud, Ref. Flower et al. (2007) found that provided the best overall agreement with the observations. Using this cloud would improve our limits by a factor on the DM annihilation/decay rate. However, the lack of a robust estimation of the uncertainties of this measurement makes us choose L1551 as reference case. Another interesting example is Barnard 68, which was found to be compatible with an ionization rate of Maret and Bergin (2007). Unfortunately, there is no precise estimation of the uncertainty of these measurements that allow us to use them reliably, but it must be noted that using these clouds would lead to at least a factor of a few stronger DM constraints than L1551. Lower values of ionization can be found in the local environment, but we leave a dedicated survey for a future work.
Interestingly, similar or even lower values of ionization have been found in infrared dark clouds (IRDCs), which are very cold (T K) and dense clouds, and can be ideal targets for DM searches. Among known IRDCs, one that has been very well studied recently is G28.37+00.07, often referred to as the DRAGON cloud, which is located around kpc away from the GC (in the so-called -kpc ring) Simon et al. (2006). This cloud has a size around L pc, a density of n cm-3 and a column density N cm-2. Ref. Entekhabi et al. (2022) studied the ionization rate and other gas properties across different sub-regions of this cloud, finding that the ionization rate in some sub-regions of the cloud was exceptionally low. Indeed, the authors of this paper chose some of these sub-regions because they were expected to gather no star formation activity. Since the ionization from photons produced by DM is expected to be uniform throughout the cloud, due to negligible variations in DM density on scales of a few parsecs, constraints on DM-induced ionization can be derived from measurements in individual zones within the cloud. In particular the region P8 of that study was found to have ionization values between s-1 and s-1 in their fiducial model. We set an upper limit on the ionization of this cloud by taking twice the highest value found in this analysis, i.e. s-1, which corresponds to our fiducial constraint. The authors of the same paper also performed an additional analysis that included more ionized species, noting that these species have more uncertain astrochemistry and are less directly linked to the induced ionization rates, which may introduce additional uncertainty into their analysis. This analysis yielded ionization rates of and . Accordingly, we present the results for the DRAGON cloud as a band, adopting constraints in the range to The P8 region has a size of pc, with a gas density around cm-3.
While the modeling of DM ionization in local clouds would not require to assume any DM distribution, we also select a cloud located quite close to the GC to demonstrate the potential of such dark clouds to probe sub-keV DM. The advantage is the higher DM density in this region, with the disadvantage of introducing uncertainties related to the DM distribution profile. For concreteness, we choose G1.4-1.8+87, with a size of around pc, and located around pc away from GC, above the Galactic plane Bhoonah et al. (2018). This cloud was found to be compatible with an ionization rate as low as Bhoonah et al. (2019), however, Ref. Farrar et al. (2020) casted serious doubts on the gas temperature used in this analysis. We will use this cloud as an example of the optimistic case of MC with low ionization rate located close to the GC. However, this should not be interpreted as a strict constraint, but rather as a forecast to illustrate the potential of similar clouds once more robust analyses are performed. To derive this limit, we choose a value of ionization of and will assume a prototypical NFW DM profile, which is the standard assumption in this kind of searches. We will call this an optimistic constraint, that will be shown in a different way than the fiducial and conservative cases, for clarity. As in the case of local MCs, more research is needed to identify more clouds close to the GC.
Ionization of UV/X-ray photons from keV DM — The ionization rate , for a MC located at x caused by photons with energy , can be calculated as Padovani et al. (2009); Phan et al. (2023)
where is the photon flux in a given location of the sky, is the photoionization cross section Padovani et al. (2009) and is the number of secondary ionizations per primary ionization Yan et al. (2001); Dogiel et al. (2014)
Here, eV is the ionization potential of H2, is the energy of the electron produced from the photoionization induced by an incoming photon with energy and eV is the average energy loss of a non-thermal electron per ionization of an H2 molecule in the interstellar medium Dalgarno et al. (1999). Precisely, the values of W we taken from Fig. 6 of Ref. Dalgarno et al. (1999). The integral in Eq. 1 extends from a minimum energy of to the maximum photon energy produced in the DM annihilation/decay, i.e. the mass of the DM particle, mχ, in the case of annihilating DM and mχ/2 in the case of decaying DM. Here, we are neglecting the Compton ionization process, which is important only above tens of keV Yan et al. (2001). This simplification is motivated by the fact that at such energies, our limits are not competitive with previous constraints.
The photon flux at a given position can be expressed as
where the factor accounts for the production of two photons, LMC is of the size of the MCs (LMC) Gabici (2022) calculated for a constant average n and is the source term for the injection of photons
Here, and are the DM thermally averaged annihilation cross section and the DM decay rate, respectively. The DM energy density at the position x is and is the -ray yield per DM interaction, that is a Dirac delta peaking at mχ, in the case of annihilation, or at mχ/2 in the case of decay.
For the local cloud L1551, we set the DM density to be the local DM density, which is GeV/cm3 Benito et al. (2021) at the Earth position kpc from the GC Evans et al. (2019). We stress that this makes our estimation independent on the DM density profile, which is one of the main uncertainties in DM indirect searches. For the case of the DRAGON cloud, which corresponds to our fiducial constraint, we assume the NFW profile Navarro et al. (1996), as it is the most usual assumption used for DM searches. We also adopt a NFW profile for our optimistic forecast with G1.4-1.8+87. We note that the local DM density value that we have taken is in the lower bound given in Ref. Benito et al. (2021), where the estimated local DM density is GeV/cm3 at 2 level.
The previous estimate assumes complete absorption of the UV photons emitted by DM within the cloud; however, this assumption does not always hold. We thus compute the opacity of the cloud () to photons emitted by DM. The resulting absorption probability, , is shown in Fig. S2 of the Supp. Material as a function of photon energy. In our calculations, specifically in Eq. 1, we multiply the integrand by this absorption probability to account for photons that escape the cloud, which suppresses our constraints towards high photon energies.
Results — In Fig. 2, we report our model-independent constraints on DM two-photon decay (left panel) and annihilation (right panel). In this energy range, the strongest constraints are from CMB spectral distortions Bolliet et al. (2021b) and heating of gas-reached dwarf galaxies (particularly, the Leo T dwarf) Wadekar and Wang (2022). Here, we show that even in the most pessimistic case of a local MC such as L1551, with a low DM density, the constraints at tens of eV are competitive with the existent bounds Wadekar and Wang (2022); Bolliet et al. (2021b); Audren et al. (2014); Poulin et al. (2016). Notably, our fiducial bound (the DRAGON cloud), becomes stronger than CMB constraints in the mass range from eV to several hundreds eV, and it is also stronger than the LeoT gas heating constraints up to a hundred eV. As commented above, the assumption of a NFW profile for the DRAGON cloud makes our constraint only a factor of stronger than assuming only the local DM density at the cloud position, and around a factor of for the annihilation rate constraint. Taking a Burkert profile Burkert (1995), which is often taken as the most pessimistic scenario in Galactic DM searches, these limits are expected to be reduced by a factor for decay and for annihilation (see Fig. S2 in section S.I. of the Supp. Material, where the parameters characterizing the DM density distribution used are given). We remind the reader, that even cuspier profiles than NFW, like the ones favored by the GC excess Ackermann and others (2017) are equally likely, and those would result in an improvement of this limit. On top of this, we note that the presence of DM minihalos or clumps within these clouds will substantially improve the predicted ionization rates, therefore leading to much stronger constraints. We remain conservative by only considering the subjacent DM distribution. Notably, for the optimistic case with G1.4-1.8+87 we get a constraint that is far stronger than previous observations below a hundred eV above the keV. We tested another interesting candidate, the cloud G357.8-4.7-55, which is around kpc Bhoonah et al. (2018) far from the GC and is also compatible with low ionization rates. Assuming NFW profile, we get a lifetime constraint one order of magnitude stronger than the L1551 MC for decaying DM.
We have also estimated what could be the effect of adding the ionization from CRs in L1551. Since CRs are expected to be the ionization agents in the interstellar medium, their effect will be dominant. In fact, previous predictions expect the ionization of L1551 to be totally explained by CRs Phan et al. (2023). We adopt the CR ionization calculated in Ref. Phan et al. (2023), pointing towards an ionization rate between s-1 and s-1 in the diffusive regime. We adopt the most conservative value of this estimation, s-1 in our calculation, and find that adding this contribution leads to around a factor of better constraints below keV for both decay and annihilation. This rough approach illustrates the improvement expected in the derived limits when accounting for the background ionization of CRs. A similar estimation could be made for clouds that are located farther away from the Solar System, but this would require a precise modeling of the distribution of low-energy CRs, and the uncertainties would be very important.
These model-independent constraints for DM producing two photons can be directly expressed as constraints in the coupling of different feebly interacting particles, such as ALPs, scalars or sterile neutrinos. We calculate our constraint on the coupling of ALPs with photons as Carenza et al. (2025b)
where is the model independent limit on the DM lifetime and ma is the ALP mass.
In Fig. 1 we report bounds for the ALPs-photon coupling, gaγ, for keV mass ALPs, compared to previous constraints. Even in the most pessimistic case of L1551, these constraints are competitive with constraints from reionization Cadamuro and Redondo (2012) and extend to higher masses. For our fiducial constraint, from the DRAGON cloud, we significantly improve on the CMB and LeoT constraints up to several hundreds of eV. Moreover, our optimistic forecast for G1.4-1.8+87, that represents what we expect from a low ionization cloud near the GC, illustrates that MC observations could be the a significant constraints in the sub-keV range for ALPs and similar particles. We also provide constraints for other well-motivated DM models in the Supplementary Material. This Supp. Material contains the following references: Yan et al. (2001); Burkert (1995); Moore and others (1999); Cirelli et al. (2011); Di Mauro (2021); Dev et al. (2021); Nguyen et al. (2024); Bolliet et al. (2021a); Wadekar and Farrar (2021); Audren et al. (2014); Poulin et al. (2016); Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021); Slatyer (2016); Fiorillo et al. (2025); Bottaro et al. (2023); Bolliet et al. (2021b); Boyarsky et al. (2019); Wadekar and Wang (2022); Bezrukov et al. (2010); Bringmann et al. (2023); Tremaine and Gunn (1979); Dev et al. (2025); Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b); Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021); Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024); Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b); Carenza et al. (2025a); Aprile and others (2019)
Discussion and conclusions — In this Letter, we investigate the ionization of neutral gas produced in dense MCs by decaying/annihilating DM. In particular, sub-keV DM could generate UV/X-ray photons that are able to produce a high and homogeneous ionization across the cloud since they are absorbed efficiently within such systems. In fact, we show that the ionization produced by the photon emission from sub-keV DM saturates the ionization that is observed in different MCs for decay lifetimes, or annihilation rates, that are not ruled out by any other constraint. Therefore, we test for the first time MCs as a target for indirect DM searches, demonstrating that they allow us to improve current constraints for decaying DM particles such as ALPs, dark photons, scalar particles or sterile neutrinos for masses from a few tens of eV to slightly below keV.
We begin by identifying local MCs as a particularly robust probe of this kind of DM. This is because estimating their ionization rate requires only knowledge of the local DM density without the need to assume a specific DM distribution model. Using the local MC L1551 we derive constraints that are similar, or even better, than the previous leading constraints. Then, we examine the case of G28.37+00.07, or the DRAGON cloud, which has been studied previously in detail and shows regions of it that are exceptionally quiescent in terms of their star formation activity, with ionization rates compatible with s-1. This constitutes what we call our fiducial constraint, and allows us improve previous constraints from CMB and LeoT gas heating up to several hundreds of eV. Finally, we use the case of G1.4-1.8+87, an MC pc away from the GC, to show the potential of using clouds near the GC. For this case, we take the values obtained in Ref. Bhoonah et al. (2019) to illustrate that, assuming an NFW DM density profile, we could reach the strongest DM bounds to date in a mass range from a few tens of eV to the keV. However, this must be taken as an optimistic forecast since the characterization of the temperature of that could was put in doubt in Ref. Farrar et al. (2020). Clouds that are closer to the GC can be found, therefore providing even stronger DM bounds. Finally, we highlight that the most promising targets for indirectly searching for sub-keV DM decays or annihilations into photons are IRDCs located near the GC but slightly above the Galactic plane. These regions are expected to be less affected by CR ionization, which is sourced in the Galactic disk. As a result, such clouds are likely to exhibit very low ionization levels, making any DM-induced effects more distinguishable. Although a more comprehensive survey of MCs will further refine the constraints presented here, our results already demonstrate the strong potential of this approach for probing different types of light DM particles.
However, even though these results are quite promising, one must keep in mind that the uncertainties in the measured ionization rates of Galactic molecular clouds are significant. These analyses require a combination of observations of molecular abundances and column densities to inform chemical models and infer the underlying ionization rates. Several factors contribute to these uncertainties. The chemistry of molecular clouds is highly sensitive to local conditions such as density, temperature, and radiation fields, which can vary across the cloud and are not always precisely known. Observational limitations, including instrumental sensitivity, line-of-sight integration, and assumptions about cloud geometry, further add to the uncertainty. Additionally, many studies report only best-fit values without fully quantified statistical errors, so the derived ionization rates often represent estimates rather than precise measurements. Despite these limitations, these analyses remain the most direct method to probe ionization processes in molecular clouds, and careful modeling allows researchers to account for uncertainties by considering conservative scenarios or quoting results as ranges. In this line, the James Webb Space Telescope has performed the first direct measurement of the ionization rate of Barnard 68 Bialy et al. (2025), which opens a window for refined ionization rate measurements in the next years.
We list a few expected future refinements below:
i) A more exhaustive survey of the measurements of dark MCs, dedicated to identify clouds with lower ionization fractions and closer to the GC. Such dedicated exploration and analysis could yield improvements of up to a few orders of magnitude in the derived DM constraints.
We show that dark clouds constitute a very interesting target for DM searches, hoping to increase the interest towards targeting dark MCs for observations.
ii) A recent paper Obolentseva et al. (2024) showed that previous measurements of the ionization rate from several dense MCs are biased, due to an inaccurate evaluation of their density. Using dust extinction maps to evaluate their density leads to an ionization rate that is lower by an average factor of . Analyzing interesting targets using this procedure could lead to constraints that are an order of magnitude better than the current ones.
iii) Probably the most obvious improvement would be that of modeling the ionization produced by CRs. Provided that CRs are expected to be the main ionization agent of dense MCs, including their contribution must certainly improve our constraints. Unfortunately, this estimation is not simple and requires an accurate knowledge of the distribution of CRs in different regions of the Galaxy and at very low energies, where our models are not robust enough to produce a reliable constraint. Reasonable estimates for local MCs already yield order-of-few improvements under conservative assumptions, as discussed above.
In conclusion, we have shown the potential impact of sub-keV DM decaying/annihilating into photons in dense gas regions, which can be specially relevant for probing ALPs and other DM particles. Our results demonstrate that astrochemical observations in MCs can serve as a novel and powerful probe for indirect searches of sub-keV DM. This approach offers particularly robust constraints across a range of particle physics models and has the potential to surpass existing bounds for DM masses between approximately eV and keV.
Acknowledgements — We thank Torsten Bringmann for fruitful discussion and Rebecca Leane for her useful comments and feedback. PDL has been supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR, and is currently supported by Ramón y Cajal RYC2024-048445-I grant, which is funded by MCIU/AEI/10.13039/501100011033 and FSE+. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the National Academic Infrastructure for Supercomputing in Sweden (NAISS) under project NAISS 2024/5-666. PC and TTQN are supported by the Swedish Research Council under contract 2022-04283. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Supplementary Material
Sub-keV dark matter can strongly ionize molecular clouds
Pedro De la Torre Luque Pierluca Carenza
Thong T. Q. Nguyen
SI Additional details
In this section, we provide additional figures that allow us to describe better our calculations. In the left panel of Fig. S1 we show the quantity J for several DM masses in a local MC like L1551. As expected, the lower the DM mass, the higher is J and, thus, the ionization rate. This behavior is due to the higher DM number density and the higher photoionization cross sections, which peaks around eV Yan et al. (2001). This can be also seen from the right panel of Fig. S1, where we show the H2 ionization rate caused by annihilating (dot-dashed line) and decaying (solid lines) DM in a local MC. The purple dashed line shows the level of ionization observed for L1551. We note that in this work we will constrain the annihilation rate into photons defined as , where is the total DM annihilation rate and is the branching ratio for photons in the final state. Similarly, we will always refer to DM decay into photons.
Another important factor in our study is the opacity of each cloud, which characterizes how efficiently photons can escape from the MC. The opacity is computed as . A very high opacity (i.e. an absorption probability close to unity) implies that photons cannot escape the molecular cloud and instead deposit all of their energy into the cloud in the form of ionization. As we see from Fig S2, while the high density of the P8 region in the DRAGON cloud makes it effectively opaque to photons with energies below 1 keV, in the L1151 and G1.4–1.8+87 clouds there exists an energy above which photons escape efficiently, producing little ionization above a few hundred keV.
Then, we show the uncertainties in our model-independent constraints on and for the DRAGON and G1.4-1.8+87 clouds associated to the assumed DM density profile. In particular, we take two extreme assumptions for the DM profile: the most pessimistic choice, which is the Burkert profile Burkert (1995) (Eq. 8) and the most optimistic choice, a Moore-like profile Moore and others (1999) (characterized by a contraction index ) (Eq. 7). Their parametric forms are the following:
For the Burkert profile (Eq. 8), we use a scale radius of kpc Cirelli et al. (2011), while for the Moore profile, we employ kpc. For the NFW profile, that is used as reference, kpc Di Mauro (2021). In all these cases the density normalization, is found by imposing that the local DM density is GeV/cm3. As we see in Fig. S3, the constraint obtained from the ionization in the DRAGON cloud is not very affected by the choice of DM density distribution, either for decay of annihilation, which makes this cloud an optimal target for robust DM searches. However, the projected constraint based on G1.4-1.8+87 varies significantly depending on the choice of DM distribution, making it less robust, but with tremendous potential under the assumption of a cuspy DM density distribution in the Galaxy.
SII Other particle models
While the case of ALPs is probably the most interesting among current light DM candidates, in this appendix we explore the constraints for other particle models coupling to photons. First, we report the constraint for scalar particles as DM, that have mixing angle with the SM Higgs Dev et al. (2021); Nguyen et al. (2024). We calculate this mixing angle from the scalar dark matter decay to two photons lifetime as
where is the fine structure constant, GeV is the electroweak vacuum expectation value and mS is the mass of the scalar particle. A comparison of our constraints with previous ones, for scalar DM, is shown in the left panel of Fig. S4. As for the case of ALPs, we are able to improve on existent constraints, such as those from CMB spectral distortions Bolliet et al. (2021a), LeoT gas heating Wadekar and Farrar (2021) or the model independent constraint on decaying DM (dDM) derived in Refs. Audren et al. (2014); Poulin et al. (2016), and our results are complementary with X-ray constraints Gewering-Peine et al. (2017); Essig et al. (2013); Foster et al. (2021), constraints from the CMB anisotropies Slatyer (2016) and stellar constraints Fiorillo et al. (2025); Bottaro et al. (2023). Interestingly, MCs can probe a region of the parameter space that cannot be currently tested from arguments of the stars dynamics.
We also compute our limit for sterile neutrinos, which have a decay of the form . The expression used for our calculation is Boyarsky et al. (2019); Bezrukov et al. (2010)
where m is the sterile neutrino mass and GF is the Fermi constant that represents the strength of the weak force. In this case, instead of producing two photons in the decay, only one is produced. Therefore, the value adopted is a factor of 2 lower than the two-photon final state. These bounds are shown in the right panel of Fig. S4, compared to previous ones. As one can see, the parameter space that is covered by MCs is already well tested by other mechanisms, specially the Lyman-alpha forest Bringmann et al. (2023). Also the Tremaine-Gunn constraint Tremaine and Gunn (1979) rules out sterile neutrinos with masses below keV. We report the optimistic constraint derived from G1.4-1.8+87, which may become relevant again if model-dependent approaches reopen this region of parameter space Dev et al. (2025).
Finally, we have also calculated these constraints for the case of vector bosons decaying into three photons. This can be the case of dark photons or the B-L vector, which are popular candidates for DM particles. In this case, the spectrum of photons is given by a continuum spectrum, instead of a line-like spectrum. The photon flux at a given position can be expressed as as Pospelov et al. (2008); Linden et al. (2024a); McDermott et al. (2018); Cheung et al. (2025); Nguyen (2025); Redondo and Postma (2009); Krnjaic et al. (2023); Linden et al. (2024b)
where . Using this spectrum, we compute Jχ→3γ for three-photon final state. The result is shown in Fig. S5, upper panel. A trend similar to the monochromatic injection appears, with a larger ionization at lower DM masses. In Fig. S5 (lower panel) we show the limit on the decay lifetime obtained for this injection of particles, for the case of L1551 and G1.4-1.8+87.
We derive constraints on well-motivated vector dark matter models, such as the dark photon and the vector boson, using decay rate limits of the vector-to-three-photon final state. Both models arise from extensions of the Standard Model by anomaly-free gauge symmetries. The dark photon interacts with Standard Model particles via kinetic mixing with the photon field Holdom (1986); Fayet (1980, 1981); Kahn et al. (2017); Fabbrichesi et al. (2020); Caputo et al. (2021), while the vector boson couples directly to the Standard Model through a flavor-universal coupling Chun and Yun (2022); Carena et al. (2004); Nguyen and Tait (2023); Smith and Tait (2024). These symmetry extensions are extensively studied in the contexts of string theory, string compactification, and vector dark matter production via quantum fluctuations during inflation Servant and Tait (2003); Graham et al. (2016); Hebecker et al. (2024); Cyncynates and Weiner (2025a, b); Carenza et al. (2025a).
Figure S6 presents the optimistic G1.4-1.8+87 constraints on vector dark matter models. We compute the decay width using the Euler-Heisenberg (EH) approximation, which is valid in the keV mass range McDermott et al. (2018); Linden et al. (2024b). The left panel shows the constraint on kinetic mixing , compared to limits from the universe’s age (4.6 Gyr), x-ray observations by INTEGRAL Linden et al. (2024b), and direct detection bounds from XENON Aprile and others (2019). For vector dark matter, we derive constraints on the flavor-universal coupling , assuming the decay proceeds dominantly via electron loops under the EH approximation. We compare this result with the INTEGRAL limit on the same three-photon final state Linden et al. (2024b) and with theoretical constraints based on neutrino decay channels and the universe’s lifetime Chun and Yun (2022).
For the dark photon, the G1.4-1.8+87 constraint surpasses the theoretical bound, as the three-photon decay dominates in this mass range. However, this region is already excluded by direct detection experiments that consider dark photon absorption. In contrast, our constraints lie above the bound set by the universe’s lifetime, assuming decay into neutrino pairs—the dominant channel for the vector boson. This behavior extends to other lepton-flavor-dependent vector dark matter models such as . Overall, our results represent conservative limits on these classes of vector dark matter models.